OpenStudy (anonymous):

(y/1+4x) (8 - x/y) (1+4x/y-x) Perform Indicated Operations

OpenStudy (luigi0210):

What are the operations?

OpenStudy (anonymous):

Well my book says multiply them all Numerator and denominator. Then cancel out like terms I'm stumped.

OpenStudy (luigi0210):

Well just multiply everything, but don't distribute them

OpenStudy (anonymous):

Yes I know but I need to find LCD and I don't know what that is for the denominator

OpenStudy (luigi0210):

When you multiply you don't really need one, it's like when you're dealing with fractions: \[\frac{1}{2}*\frac{5}{3}=\frac{5}{6}\] It'd be like this right?

OpenStudy (anonymous):

Yes. But then I need to cancel out like terms and sometimes that is where i find it confusing because I will cancel out 4xy with a y because i see a y. But these two terms are not the same

OpenStudy (luigi0210):

Here I'll get you started: \[\frac{y}{1+4x}*\frac{8-x}{y}*\frac{1+4x}{y-x}=\frac{(y)(8-x)(1+4x)}{(1+4x)(y)(y-x)}\] Now start canceling :)

OpenStudy (anonymous):

do all three y's get canceled out?

OpenStudy (anonymous):

my answer is 8x^2?

OpenStudy (anonymous):

8-x^2 i mean

OpenStudy (luigi0210):

y and (1+4x) cancel, how did you get an x^2?

OpenStudy (anonymous):

or perhaps the answer is just 8 lol

OpenStudy (anonymous):

8 - x/y is in one parethesis

OpenStudy (anonymous):

i canceled out both x's and i am left with 8

OpenStudy (luigi0210):